Name: John Macomber

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. (10) Transitions between paragraphs are logical – 10

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 15

Total: 100 (A)

Style and English Usage – The paper

1. (20) Is clear, direct, and appropriate in tone – 20

2. (10) Is free of misspellings – 9

3. (15) Is free of grammatical mistakes – 15

4. (10) Avoids incomplete sentences – 10

5. (10) Avoids run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 14

8. (10) Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments: John – This is a good, convincing argument that it was an abuse of power that lead to some of the questionable aspects of the treatment of Henrietta Lacks. The analogy with the recent financial crisis is interesting and well presented. There are a still a few places where things could be tightened up and polished in the writing, but that is mostly very good too.